OpenStudy (anonymous):

Which is the set of roots for the equation? x2 + 3x - 18 = 0 A. {18,-1} B. {-6,3} C. {-2,9} D. {9} Which is the set of roots for the equation? x2 + 3x - 18 = 0 A. {18,-1} B. {-6,3} C. {-2,9} D. {9} @Mathematics

OpenStudy (anonymous):

B

OpenStudy (anonymous):

yup b is correct

OpenStudy (anonymous):

Use the quadratic formula: \[(-b \pm \sqrt{b^{2}-4ac})\div2a\] To find the roots for a quadratic, to the form of \[ax^{2} + bx + c\] In this case, a = 1, b = 3 and c = -18 thus: \[(-3\pm \sqrt{3^{2}-4(1\times18)})\div(2\times1)\] \[=(-3\pm \sqrt{9-(-72)})\div2\] \[=(-3\pm9)\div2\]

OpenStudy (anonymous):

Another way to look at this problem is by generating factors. -18 = -(6 * 3). To get +3x, you need 6 to be positive and 3 to be negative. So, the factors are (x+6)(x-3). This equation is zero if x = -6 or x = 3.